Assume that there are 4 firmssupplying a homogenous product .they habe identical cost function given by c(q)=40 q .if demand curve fr the industry is given by u =100_q find the equilibrium industry output if prodicers are cournot competitors.what would be the resultant market price ? What are the profit of each firm
1
Expert's answer
2017-06-05T08:48:10-0400
TR=(100-Q)*Q=100Q- 2Q^2

MR=100-2Q

100-2q=40Q

42Q=100

Q=2.38

P=97.62

Profit=(2.38*97.62-40*2.38)/4=32.84